your montly mortgage payment ( principal plus interest) is $1,500.00. If you have a 30 years loan with a fixed interest rate of 6% compounded monthly, how much did you borrow?
Answer #1

We know, A = P (1 + r/m)mt, where A = 1500, r = 6% m =1, t =30

1500 = P (1 + 0.06)30

P = 1500 /(1.06)30

P = 1500 / 5.74 = 261.3

So 261.3 amount was borrowed initally.